OpenStudy (anonymous):

Mark and Jose have a combined weight of 270 pounds. Mark weighs 30 pounds less than twice Jose's weight. How much do Mark and Jose weigh?

OpenStudy (anonymous):

Set equations: let mark's weight = x and jose's weight = y then : x + y = 270 x = 2y - 30 do yo see where i got the equations?

OpenStudy (anonymous):

good.

OpenStudy (campbell_st):

write it in terms of 1 variable let jose weight = x then mark's weight = 2x - 30 their combined weight is x + 2x - 30 = 270 so 3x - 30 = 270 solve for x add 30 to both sides 3x = 300 divide by 3 x = 100 Jose = 100 lbs Mark = 170 lbs

OpenStudy (anonymous):

that helps thanks~